# This version: August 29, 2025 )

### v3

@@ 4 Identification with Strategic Reports
This section studies identification of causal effects when students are strategic in reporting their rankings of acceptable schools. Strategic reports make $P_{j}$ generally different from $Q_{j}$, and $Q_{j}$ is not observed. In contrast to Proposition 1, controlling for $P_{j}$ does not identify the parameters of interest in (3), which condition on $Q_{j}$. Moreover, controlling for $P_{j}$ may break the internal validity of the RD in some cases because $P_{j}$ is a variable subject to manipulation by agents. The second reason is more subtle and we explain it in Section 4.2 below in terms of our recurring SD Example (Part III).
We propose a two-step identification approach that is robust to strategic reporting. In the first step, the researcher characterizes the set of true local preferences $Q_{j}$ that is compatible with the data and appropriate behavioral assumptions. In the second step, the researcher controls for the constructed local preference sets and partially identifies the parameters in (3). We discuss the first and second steps in Sections 4.1 and 4.2, respectively. Note that our two-step approach differs from the usual two-step control function approach in econometrics. The usual approach is to point-identify the control variable in the first step, while our approach involves partially identifying the control variable. Thus, we refer to our two-step procedure as a control mapping approach.
Section 4.1 presents several tools for the identification of local preference sets. These tools rely on assumptions known to be appropriate in SD and DA contexts, although we do not rule out their applicability in contexts with other mechanisms; e.g., the TTC mechanism satisfies one of our assumptions, such that some of the tools from Section 4.1 are still useful. More generally, researchers may utilize preference identification tools that work under alternative assumptions, for example, the methods of Agarwal and Somaini (2018) and Fack et al. (2019). Either way, the researcher must construct a set of local preferences for each individual in the first step.
Section 4.2 describes the second step of our procedure. This step features high-level assumptions imposed on the local preference sets such that the researcher is not restricted to the methods of Section 4.1.
@@ 4.1 Partial Identification of Local Preferences
This section provides tools for set identification of local preferences using assumptions on agents’ behavior and the mechanism. These assumptions are specific to this subsection, and we motivate them with reference to the context of the constrained DA mechanism studied by Haeringer and Klijn (2009). Haeringer and Klijn (2009) study a game where students submit constrained preference rankings and a mechanism matches students to schools as a function of $P$, $\bm{\mathrm{R}}$, and schools’ capacities. Although the unconstrained DA mechanism is strategy proof, many real-world implementations of DA restrict the number of schools that students can submit in their rankings. In this case, there is a cap $K<J$ such that $1\leq|P|\leq K$, and the submitted ranking $P$ is generally different from the list of acceptable schools in $Q$. When implemented in this way, the DA mechanism is not strategy-proof, and there are no dominant strategies. Strategyproofness also breaks down if, instead of facing a cap, students incur an application cost as a function of the number of schools submitted (Fack et al., 2019).
Lemma 4.2 by Haeringer and Klijn (2008) shows that, if a mechanism is strategy proof when $K=J$, then, in the game with $K<J$, any constrained ranking of schools is weakly dominated by the same set of schools ranked according to true preferences. This result implies that a student cannot lose and may possibly gain by taking any arbitrary list with less than or equal to $K$ schools, dropping the unacceptable schools, and ranking the acceptable schools according to her true preferences. A further implication is that if a student’s number of acceptable schools is less than or equal to $K$, then her dominant strategy is to submit her true list of acceptable schools (Proposition 4.2 of Haeringer and Klijn (2009)). These implications give rise to a class of undominated strategies according to the following definitions of partial order.
@@
Lemma 2 in Section A.2 of the appendix summarizes the implications of the result on partial orders from Haeringer and Klijn (2008, 2009) in terms of our Definition 6. In short, for a student with true preferences $Q$, any $P$ is weakly dominated by a weak partial order $P^{*}$ of $Q$ that has the same acceptable schools as $P$; in turn, $P^{*}$ is weakly dominated by a strong partial order $P^{**}$ of $Q$ that contains the same set of acceptable schools as $P^{*}$. Every strong partial order is a weak partial order, but the converse is not true. Our definition of a weak partial order strategy is similar to the definition of the dropping strategy from Kojima and Pathak (2009).
Assuming that agents always submit a strong partial order implies they reveal their true ordered list of acceptable schools whenever they submit $P$ with fewer schools than the cap $K$. This could be a strong behavioral assumption in some contexts where agents have more than $K$ acceptable schools but have a strong expectation that they will gain admission to a smaller-than-$K$ set of schools. In this case, they may submit $|P|<K$ not because it reflects their full list of acceptable schools but simply because they may not want to incur the all the costs of ranking all schools up to $K$. In the rest of this subsection, we consider mechanisms that impose a cap $K$ on $P$ and assume that students submit a weak partial order of their true preferences.
@@
An assumption maintained throughout this paper is that the cutoff characterization from Definition 2 applies to the mechanism. This assumption says that $\mu(\omega)=Q(B(\bm{\mathrm{S}}(\omega))$ for every $\omega\in\Omega$. Azevedo and Leshno (2016) show that stability is equivalent to cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and cutoffs that equal the minimum score of the admitted students in each school. Thus, it is worth discussing the stability of the constrained DA mechanism. Theorem 6.3 from Haeringer and Klijn (2009) demonstrates that any Nash equilibrium in constrained DA where $\bm{\mathrm{R}}$ satisfies Ergin acyclicity leads to a stable matching in the finite economy.999 Ergin acyclicity ensures that no student can block a potential improvement for any two other students without affecting her own assignment. See Ergin (2002) for the formal definition. Even without Ergin acyclicity, some Nash equilibria still produce stability. SD always satisfies Ergin acyclicity, and thus every Nash equilibrium in SD produces a stable matching. Fack et al. (2019) also study the constrained DA mechanism. They extend Theorem 6.3 from Haeringer and Klijn (2009) to (pure-strategy) Bayesian Nash equilibria in the continuum economy (Proposition A3, Online Appendix A.2.5, Fack et al. (2019)). Fack et al. (2019) also provide primitive conditions for finite economies where students play partial orders to converge to a continuum economy with a stable equilibrium (Proposition 5, Fack et al. (2019)). They further provide a test for implications of stability and find no empirical or simulation evidence against it. In the context of the constrained TTC mechanism, any Nash equilibrium leads to a stable matching as long as $\bm{\mathrm{R}}$ satisfies Kesten acyclicity (Theorem 6.4 from Haeringer and Klijn (2009)). Therefore, constrained SD, DA, and TTC all satisfy the assumption of cutoff characterization as in Definition 2 with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ under the appropriate conditions.
@@
The DA mechanism, constrained or unconstrained, produces a matching that is stable wrt reported preferences $P$. Stability wrt $P$ leads to a cutoff characterization wrt $P$ according to the work of Azevedo and Leshno (2016). This cutoff characterization has scores $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs that equal the smallest scores of admitted students in each school. In other words, this is the same cutoff characterization from Definition 2 except that $Q$ is replaced with $P$. Cutoff characterization wrt $P$ is natural in DA but not necessarily in other mechanisms, so we state it in the following assumption.
@@
This example illustrates why an RD at $c_{4}$ that controls for $P_{4}=(4,2)$ might be problematic. The range of possibilities for true preference types is different between individuals above and below the cutoff. We see types $(4,1)$, $(4,2)$, and $(4,3)$ above the cutoff but only type $(4,2)$ below the cutoff. Suppose in an extreme case there is a nonzero fraction of individuals with types $(4,1)$ or $(4,3)$ right above the cutoff. True preferences arguably affect outcomes and this implies that the distribution of outcomes conditional on $P_{4}=(4,2)$ and $S_{1}=s$ changes as $s$ crosses the cutoff $c_{4}$ even without a treatment effect. This extreme case breaks internal validity of RD identification. Even if the fraction of types $(4,1)$ or $(4,3)$ at the cutoff is zero and RD identification is valid, a continuous but sudden increase in the fraction of such agents on the right of the cutoff may bring severe bias issues and make inference impossible (e.g., Bertanha and Moreira (2020), in particular, Sections 5 and A.6).
For the issue to arise when we control for $P_{4}=(4,2)$, there must be at least a fraction of agents with true local preferences $(4,1)$ and $(4,3)$ who change their submission behavior discontinuously as a function of $S_{1}$ as $S_{1}$ crosses the value of $c_{4}$; i.e., they must change $P$ such that $P_{4}\neq(4,2)$ below the cutoff and $P_{4}=(4,2)$ above the cutoff. Intuitively, this requires these agents to have good ex-ante knowledge about the ex-post value of the cutoff $c_{4}$. The more knowledge they have about the ex-post cutoff values, the sharper will be their change in behavior around the cutoff, and the closer we are to the extreme case mentioned above. Section A.5 in the appendix presents a numerical example of an economy with agents that maximize expected utility in face of cutoff uncertainty and exhibit such discontinuous behavior at the ex-post cutoffs. Finally, Section 5.1 displays empirical evidence from Chile that supports agents having good a priori knowledge of cutoffs to the point of influencing their submissions.
@@
Agarwal and Somaini (2018) propose a general framework to rationalize strategic reporting as the optimal solution to an expected utility maximization problem. In this framework, agents have private information about their preferences and scores and form beliefs about the distribution of other people’s preferences and scores. These beliefs plus knowledge of the mechanism lead the rational agent to derive probabilities of admission to the various schools as a function of the agent’s private information and expectations about other agents. The agent then chooses the submission $P$ that maximizes her expected utility.
@@
Assumption 5 says that students correctly anticipate which schools will be uniformly more accessible after the matching assignment. For example, agents may learn this information by observing past realizations of the matching in the economy. If a school $e$ is well known to be accessible to everyone who has access to school $d$, then it is natural for a student to expect to have access to $e$ if she ever has access to school $d$. Note that the assumption does not pin down the expected probability of admission or the set of schools to which the student will have access in the ex post economy. It restricts only the expected hierarchy of school access according to $UMAS$. This assumption has implications for the joint distribution of $(P,Q)$.
@@ 4.2 Partial Identification of Causal Effects
Recall that the construction of the random set $\bm{\mathrm{Q}}_{j}$ depends on assumptions regarding the behavior of agents when they submit $P$. For example, Section 4.1 characterizes $\bm{\mathrm{Q}}_{j}$ by assuming weak partial order and cutoff characterization wrt $P$. Alternatively, the identification approach of Agarwal and Somaini (2018) makes different types of assumptions on agents’ expectations and requires data variation in the choice environment. The theoretical credibility of these types of assumptions depends on the mechanism faced by agents; in practice, the assumptions have testable implications for what we should observe in the data. It is therefore useful to characterize a falsification test based on these implications to aid researchers in screening out assumptions rejected by the data. Such a test can rely on the fact that the right-hand sides of the inequalities in Proposition 4 must provide a lower bound for a probability mass function if the model assumptions are correct. For any partition $\mathcal{A}$ of $\mathcal{J}^{0}\times\mathcal{J}^{0}$, we must have $\sum_{A\in\mathcal{A}}\mathbb{P}\left[Q_{j}\in A|S_{j}=c_{j}\right]=1$ for any given distribution in the sharp set of Proposition 4. Thus, the same sum applied to the right-hand sides of the inequalities above must be less than or equal to one.
Corollary 2 (Model’s Falsification Test).
@@
The same idea applies for individuals just below the cutoff. Select all individuals whose $\bm{\mathrm{Q}}_{j}$ contain the comparable pair of interest, $(j,k)\in\mathcal{P}$. The fraction of those who have $Q_{j}=(j,k)$ equals
$\delta_{j,k}^{-}=\frac{\mathbb{P}\left[Q_{j}=(j,k)|S_{j}=c_{j}\right]}{\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset|S_{j}=c_{j}^{-}\right]},$
and the sharp lower bound on $\delta_{j,k}^{-}$ is
$\underline{\delta}_{j,k}^{-}=\frac{\underline{p}_{j,k}}{\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset|S_{j}=c_{j}^{-}\right]}.$
@@
The expressions for the bounds simplify significantly in the cases where $Y$ is binary or continuous. For the sake of brevity, we relegate the detailed formulas to the Appendix (Section B.7). The bounds in Proposition 5 build on the work by Horowitz and Manski (1995). To see the intuition, consider the case where $g(Y)=Y$ is a continuous random variable, and focus on individuals just above the cutoff. Among all individuals in the subpopulation with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$, a fraction ${\delta}^{+}_{j,k}$ of them has $Q_{j}=(j,k)$ and $Y=Y(j)$ by the cutoff characterization. We do not know who these individuals are among those in the subpopulation. However, the lowest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all such individuals are located at the lower tail of the distribution of outcomes in the subpopulation. Likewise, the highest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all of that same fraction of individuals are located in the upper tail of the distribution of outcomes. We do not know ${\delta}^{+}_{j,k}$, but we do know that it is no smaller than $\underline{\delta}^{+}_{j,k}>0$. The bounds only grow wider as the fraction ${\delta}^{+}_{j,k}$ decreases, so the bounds evaluated at ${\delta}^{+}_{j,k}=\underline{\delta}^{+}_{j,k}$ take into account all possible values for ${\delta}^{+}_{j,k}$.
Although intuitive and analytically simple, these bounds are not necessarily sharp because $\underline{\delta}^{+}_{j,k}$ is not exogenously given as considered by Horowitz and Manski (1995); $\underline{\delta}^{+}_{j,k}$ is constructed from the marginal distribution of $Q_{j}$ but there could be additional identifying information in the joint distribution of $Q_{j}$ and $Y$. Providing a complete characterization of the sharp bounds is complex because it involves deriving bounds on the joint distribution of potential outcomes and $Q_{j}$, which may not be practical when the potential outcomes are continuous. For the sake of simplicity, we relegate the sharp characterization to Section A.3 in the appendix. The bounds of Proposition 5 contain the sharp bounds of Section A.3 as long as our model assumptions are true. The lack of sharpness may not matter in practice when Proposition 5 yields tight bounds for a given dataset. However, if our assumptions are not true, the bounds of Proposition 5 may not contain the sharp bounds of Section A.3. That said, we may obtain tight bounds from the data using Proposition 5, but this does not mean they contain the true parameter (Kédagni et al., 2020). Therefore, it is advisable to assess the testable implications of Corollary 2 as a matter of routine.

### v4

@@ 4 Identification with Strategic Reports
This section studies identification of causal effects when students are strategic in reporting their rankings of acceptable schools. Strategic reports make $P_{j}$ generally different from $Q_{j}$, and $Q_{j}$ is not observed. In contrast to Proposition 1, controlling for $P_{j}$ likely identifies a weighted average of various average treatment effect parameters; however, the weights are unknown because they depend on the unknown mapping $\omega\mapsto Q_{j}(\omega)$. In addition, there is the possibility of a more extreme form of strategic behavior, that is, agents may change their preference submission discontinuously at the cutoff. We provide empirical evidence of such behavior in Section 5.1. In this case, controlling for $P_{j}$ may break internal validity of the RD, and the strategy of controlling for $P_{j}$ no longer identifies a weighted average of average treatment effect parameters. Although our methods are also robust to this extreme possibility, we emphasize that it is not the only situation where researchers may find our methods useful. Our methods seek to identify (3) but controlling for $P_{j}$ does not identify (3) under general forms of strategic behavior, regardless if it is discontinuous or not.
We propose a two-step identification approach. In the first step, the researcher characterizes the set of true local preferences $Q_{j}$ that is compatible with the data and appropriate behavioral assumptions. In the second step, the researcher controls for the constructed local preference sets and partially identifies the parameters in (3). We discuss the first and second steps in Sections 4.1 and 4.2, respectively. Note that our two-step approach differs from the usual two-step control function approach in econometrics. The usual approach is to point-identify the control variable in the first step, while our approach involves partially identifying the control variable. Thus, we refer to our two-step procedure as a control mapping approach.
Section 4.1 presents several tools for the identification of local preference sets. These tools rely on assumptions known to be appropriate in SD and DA contexts, although we do not rule out their applicability in contexts with other mechanisms; e.g., the TTC mechanism satisfies one of our assumptions, such that some of the tools from Section 4.1 are still useful. More generally, researchers may utilize preference identification tools that work under alternative assumptions, for example, the methods of agarwal2018 and fack2019. Either way, the researcher must construct a set of local preferences for each individual in the first step.
Section 4.2 describes the second step of our procedure. This step features high-level assumptions imposed on the local preference sets such that the researcher is not restricted to the methods of Section 4.1. In particular, depending on the data and the strength of assumptions in the first step, each individual local preference set may collapse to a unit set. This leads to point identification in our second step. For example, this will be the case if researchers utilize the parametric identification tools of agarwal2018 or fack2019 in the first step.
@@ 4.1 Partial Identification of Local Preferences
This section provides tools for set identification of local preferences using assumptions on agents’ behavior and the mechanism. These assumptions are specific to this subsection, and we motivate them with reference to the context of the constrained DA mechanism studied by haeringer2009. haeringer2009 study a game where students submit constrained preference rankings and a mechanism matches students to schools as a function of $P$, $\bm{\mathrm{R}}$, and schools’ capacities. Although the unconstrained DA mechanism is strategy proof, many real-world implementations of DA restrict the number of schools that students can submit in their rankings. In this case, there is a cap $K<J$ such that $1\leq|P|\leq K$, and the submitted ranking $P$ is generally different from the list of acceptable schools in $Q$. When implemented in this way, the DA mechanism is not strategy-proof, and there are no dominant strategies. Strategyproofness also breaks down if, instead of facing a cap, students incur an application cost as a function of the number of schools submitted (fack2019).
Lemma 4.2 by haeringer2008 shows that, if a mechanism is strategy proof when $K=J$, then, in the game with $K<J$, any constrained ranking of schools is weakly dominated by the same set of schools ranked according to true preferences. This result implies that a student cannot lose and may possibly gain by taking any arbitrary list with less than or equal to $K$ schools, dropping the unacceptable schools, and ranking the acceptable schools according to her true preferences. A further implication is that if a student’s number of acceptable schools is less than or equal to $K$, then her dominant strategy is to submit her true list of acceptable schools (Proposition 4.2 of haeringer2009). These implications give rise to a class of undominated strategies according to the following definitions of partial order.
@@
Lemma 2 in Section A.2 of the appendix summarizes the implications of the result on partial orders from haeringer2008; haeringer2009 in terms of our Definition 6. In short, for a student with true preferences $Q$, any $P$ is weakly dominated by a weak partial order $P^{*}$ of $Q$ that has the same acceptable schools as $P$; in turn, $P^{*}$ is weakly dominated by a strong partial order $P^{**}$ of $Q$ that contains the same set of acceptable schools as $P^{*}$. Every strong partial order is a weak partial order, but the converse is not true. Our definition of a weak partial order strategy is similar to the definition of the dropping strategy from kojima2009incentives.
Assuming that agents always submit a strong partial order implies they reveal their true ordered list of acceptable schools whenever they submit $P$ with fewer schools than the cap $K$. This could be a strong behavioral assumption in some contexts where agents have more than $K$ acceptable schools but have a strong expectation that they will gain admission to a smaller-than-$K$ set of schools. In this case, they may submit $|P|<K$ not because it reflects their full list of acceptable schools but simply because they may not want to incur the costs of ranking all schools up to $K$. In the rest of this subsection, we consider mechanisms that impose a cap $K$ on $P$ and assume that students submit a weak partial order of their true preferences.
@@
An assumption maintained throughout this paper is that the cutoff characterization from Definition 2 applies to the mechanism. This assumption says that $\mu(\omega)=Q(B(\bm{\mathrm{S}}(\omega))$ for every $\omega\in\Omega$. azevedo2016 show that stability is equivalent to cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and cutoffs that equal the minimum score of the admitted students in each school. Thus, it is worth discussing the stability of the constrained DA mechanism. Theorem 6.3 from haeringer2009 demonstrates that any Nash equilibrium in constrained DA where $\bm{\mathrm{R}}$ satisfies Ergin acyclicity leads to a stable matching in the finite economy. [^11] Ergin acyclicity ensures that no student can block a potential improvement for any two other students without affecting her own assignment. See ergin2002 for the formal definition. Even without Ergin acyclicity, some Nash equilibria still produce stability. SD always satisfies Ergin acyclicity, and thus every Nash equilibrium in SD produces a stable matching. fack2019 also study the constrained DA mechanism. They extend Theorem 6.3 from haeringer2009 to (pure-strategy) Bayesian Nash equilibria in the continuum economy (Proposition A3, Online Appendix A.2.5, fack2019). fack2019 also provide primitive conditions for finite economies where students play partial orders to converge to a continuum economy with a stable equilibrium (Proposition 5, fack2019). They further provide a test for implications of stability and find no empirical or simulation evidence against it. In the context of the constrained TTC mechanism, any Nash equilibrium leads to a stable matching as long as $\bm{\mathrm{R}}$ satisfies Kesten acyclicity (Theorem 6.4 from haeringer2009). Therefore, constrained SD, DA, and TTC all satisfy the assumption of cutoff characterization as in Definition 2 with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ under the appropriate conditions.
@@
The DA mechanism, constrained or unconstrained, produces a matching that is stable wrt reported preferences $P$. Stability wrt $P$ leads to a cutoff characterization wrt $P$ according to the work of azevedo2016. This cutoff characterization has scores $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs that equal the smallest scores of admitted students in each school. In other words, this is the same cutoff characterization from Definition 2 except that $Q$ is replaced with $P$. Cutoff characterization wrt $P$ is natural in DA but not necessarily in other mechanisms, so we state it in the following assumption.
@@
This example illustrates why an RD at $c_{4}$ that controls for $P_{4}=(4,2)$ might be problematic. The range of possibilities for true preference types is different between individuals above and below the cutoff. We see types $(4,1)$, $(4,2)$, and $(4,3)$ above the cutoff but only type $(4,2)$ below the cutoff. Suppose in an extreme case there is a nonzero fraction of individuals with types $(4,1)$ or $(4,3)$ right above the cutoff. True preferences arguably affect outcomes and this implies that the distribution of outcomes conditional on $P_{4}=(4,2)$ and $S_{1}=s$ changes as $s$ crosses the cutoff $c_{4}$ even without a treatment effect. This extreme case breaks internal validity of RD identification. Even if the fraction of types $(4,1)$ or $(4,3)$ at the cutoff is zero and RD identification is valid, a continuous but sudden increase in the fraction of such agents on the right of the cutoff may bring severe bias issues and make inference impossible (e.g., bertanha_impossible, in particular, Sections 5 and A.6).
For the issue to arise when we control for $P_{4}=(4,2)$, there must be at least a fraction of agents with true local preferences $(4,1)$ and $(4,3)$ who change their submission behavior discontinuously as a function of $S_{1}$ as $S_{1}$ crosses the value of $c_{4}$; i.e., they must change $P$ such that $P_{4}\neq(4,2)$ below the cutoff and $P_{4}=(4,2)$ above the cutoff. Intuitively, this requires these agents to have good ex-ante knowledge about the ex-post value of the cutoff $c_{4}$. The more knowledge they have about the ex-post cutoff values, the sharper will be their change in behavior around the cutoff, and the closer we are to the extreme case mentioned above. Section A.6 in the appendix presents a numerical example of an economy with agents that maximize expected utility in face of cutoff uncertainty and exhibit such discontinuous behavior at the ex-post cutoffs. An interesting feature of that example is that agents do not need to have perfect knowledge of the cutoffs to exhibit discontinuous behavior. Finally, Section 5.1 displays empirical evidence from Chile that supports agents having good a priori knowledge of cutoffs; we also show that submission behavior changes near ex post cutoffs and the change is discontinuous in some cases.
@@
agarwal2018 propose a general framework to rationalize strategic reporting as the optimal solution to an expected utility maximization problem. In this framework, agents have private information about their preferences and scores and form beliefs about the distribution of other people’s preferences and scores. These beliefs plus knowledge of the mechanism lead the rational agent to derive probabilities of admission to the various schools as a function of the agent’s private information and expectations about other agents. The agent then chooses the submission $P$ that maximizes her expected utility.
@@
Assumption 5 says that students correctly anticipate which schools will be uniformly more accessible after the matching assignment. For example, agents may learn this information by observing past realizations of the matching in the economy. If a school $e$ is well known to be accessible to everyone who has access to school $d$, then it is natural for a student to expect to have access to $e$ if she ever has access to school $d$. Note that the assumption does not pin down the expected probability of admission or the set of schools to which the student will have access in the ex-post economy. It restricts only the expected hierarchy of school access according to $UMAS$. This assumption has implications for the joint distribution of $(P,Q)$.
@@ 4.2 Partial Identification of Causal Effects
Recall that the construction of the random set $\bm{\mathrm{Q}}_{j}$ depends on assumptions regarding the behavior of agents when they submit $P$. For example, Section 4.1 characterizes $\bm{\mathrm{Q}}_{j}$ by assuming weak partial order and cutoff characterization wrt $P$. Alternatively, the identification approach of agarwal2018 makes different types of assumptions on agents’ expectations and requires data variation in the choice environment. The theoretical credibility of these types of assumptions depends on the mechanism faced by agents; in practice, the assumptions have testable implications for what we should observe in the data. Section A.4 in the appendix provides a testable implication of our assumptions.
@@
The same idea applies for individuals just below the cutoff:
$\delta_{j,k}^{-}=\frac{\mathbb{P}\left[Q_{j}=(j,k)|S_{j}=c_{j}\right]}{\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset|S_{j}=c_{j}^{-}\right]},\hskip 28.45274pt\underline{\delta}_{j,k}^{-}=\frac{\underline{p}_{j,k}}{\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset|S_{j}=c_{j}^{-}\right]}.$
@@
The expressions for the bounds simplify significantly in the cases where $Y$ is binary or continuous. For the sake of brevity, we relegate the detailed formulas to the Appendix (Section B.7). The bounds in Proposition 5 build on the work by horowitz1995. To see the intuition, consider the case where $g(Y)=Y$ is a continuous random variable, and focus on individuals just above the cutoff. Among all individuals in the subpopulation with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$, a fraction ${\delta}^{+}_{j,k}$ of them has $Q_{j}=(j,k)$ and $Y=Y(j)$ by the cutoff characterization. We do not know who these individuals are among those in the subpopulation. However, the lowest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all such individuals are located at the lower tail of the distribution of outcomes in the subpopulation. Likewise, the highest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all of that same fraction of individuals are located in the upper tail of the distribution of outcomes. We do not know ${\delta}^{+}_{j,k}$, but we do know that it is no smaller than $\underline{\delta}^{+}_{j,k}>0$. The bounds only grow wider as the fraction ${\delta}^{+}_{j,k}$ decreases, so the bounds evaluated at ${\delta}^{+}_{j,k}=\underline{\delta}^{+}_{j,k}$ take into account all possible values for ${\delta}^{+}_{j,k}$.
Although intuitive and analytically simple, these bounds are not necessarily sharp because $\underline{\delta}^{+}_{j,k}$ is not exogenously given as considered by horowitz1995; $\underline{\delta}^{+}_{j,k}$ is constructed from the marginal distribution of $Q_{j}$ but there could be additional identifying information in the joint distribution of $Q_{j}$ and $Y$. Providing a complete characterization of the sharp bounds is complex because it involves deriving bounds on the joint distribution of potential outcomes and $Q_{j}$, which may not be practical when the potential outcomes are continuous. For the sake of simplicity, we relegate the sharp characterization to Section A.3 in the appendix. The bounds of Proposition 5 contain the sharp bounds of Section A.3 as long as our model assumptions are true. The lack of sharpness may not matter in practice when Proposition 5 yields tight bounds for a given dataset. However, if our assumptions are not true, the bounds of Proposition 5 may not contain the sharp bounds of Section A.3. That said, we may obtain tight bounds from the data using Proposition 5, but this does not mean they contain the true parameter (kedagni2020). Therefore, it is advisable to assess the testable implications of Corollary 2 as a matter of routine.